OpenStudy (anonymous):

2. If y(x-1)=z then x=

OpenStudy (anonymous):

Start by dividing both sides by y

OpenStudy (phi):

here is the first step. Like Caleb said, divide both sides by y \[ \frac{y(x-1)}{y} = \frac{z}{y} \] the left side can be simplified, Can you see how?

OpenStudy (anonymous):

yes...thanks so much.

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

i got x=z/y+1

OpenStudy (phi):

you could also write the answer as \[ \frac{z}{y}+\frac{y}{y} \] or \[ \frac{z+y}{y} \] just in case you see that in a multiple choice answer.

OpenStudy (anonymous):

thanks very much!!! :)